# Align Attention Heads Before Merging Them: An Effective Way for Converting MHA to GQA

### v1

@@ Abstract
Large language models have been shown to perform well on a variety of natural language processing problems. However, as the model size and the input sequence’s length increase, the rapid increase of KV Cache significantly slows down inference speed. Therefore GQA model, as an alternative to MHA model, has been widely introduced into LLMs. In this work, we propose a low-cost method for pruning MHA models into GQA models with any compression ratio of key-value heads. Our method is based on $\mathit{L_{0}}$ masks to gradually remove redundant parameters. In addition, we apply orthogonal transformations to attention heads without changing the model to increase similarity between attention heads before pruning training, in order to further improve performance of the model. Our method can be compatible with rotary position embedding (RoPE), which means the model after training can be fully adapted to the mainstream standard GQA framework. Experiments demonstrate that our strategy can compress up to 87.5% of key-value heads of the LLaMA2-7B model without too much performance degradation, just achieved through supervised fine-tuning.
@@ 1 Introduction
Recently, Large language models (LLMs) Radford et al. (2018); Brown et al. (2020); Ouyang et al. (2022) show remarkable performance on a variety of natural language processing tasks. However, since most LLMs are based on Transformer architecture Vaswani (2017), the expansion of the sequence length during inference results in a linear increase in the memory footprint of KV Cache, significantly slowing down model inference. Therefore, reducing the size of KV Cache is a key issue for LLMs.
@@
Multi-query attention (MQA) Shazeer (2019) and grouped-query attention (GQA) Ainslie et al. (2023) reduce KV Cache by allowing multiple attention heads to share a single key-value head, which is simple and effective. Since GQA has better inference stability and performance, it has been widely used in LLaMA 2 Touvron et al. (2023), LLaMA 3 Dubey et al. (2024), Qwen2 Yang et al. (2024), Mistral Jiang et al. (2023) and other LLMs Liu et al. (2024); Zhang et al. (2024).
A method for converting MHA model to GQA model is also proposed in Ainslie et al. (2023). They construct each group key and value head by mean-pooling all the original heads within that group and then uptrain the model to restore the model performance. However, the computational resources required for uptraining are unaffordable in most conditions.
Figure 1: An illustration of our pruning training process. The training loss consists of two parts: distillation loss between teacher model and pruning model, and pruning loss to impose key-value projection matrices gradually transferred to new ones. After pruning, original key-value projection matrices will be discarded, then we get a standard GQA model.
In this work, we use $\mathit{L_{0}}$ masks Louizos et al. (2017) to transfer original key-value projection matrices to new ones. Figure 1 shows our concept. In addition, based on the idea of computational invariance in the model Ashkboos et al. (2024), we apply orthogonal transformations to the matrices of attention heads without changing the model before pruning training: We measure the similarity between KV Caches of different attention heads after orthogonal transformation, then divide attention heads with high similarity of KV Caches into the same group. The similarity among the KV Caches derived from the attention heads within the same group is then maximized through an orthogonal transformation. Experiments show that this method can significantly improve the performance of the pruned model after the transformation of attention heads. Yu et al. (2024) is the most relevant work to ours; however, their method is not fully compatible with RoPE Su et al. (2024). Our contributions are as follows.
@@
  • We propose a general and low-cost method for converting MHA structure to GQA, which can compress the key and value heads to any percentage and basically restore performance after supervised fine-tuning.
  • Our study provides a new perspective for evaluating the similarity between attention heads, which presents new insights for future research related to compressing KV Cache.
@@
  • We conduct experiments on pruning training the LLaMA2-7B model Touvron et al. (2023) into GQA-16, GQA-8, and GQA-4, separately. The model performance does not decrease significantly compared to that of the full-size model.
    Figure 2: In each block of LLaMA2-7B, the average cosine similarity is calculated between every two key and value caches. For convenience, the average similarities shown in this figure are their absolute values. It can be seen that most pairs of KV Caches are almost orthogonal. This may explain why directly merging key-value heads causes significant loss.
@@ 2.2 Transfer MHA to GQA
Ainslie et al. (2023) proposes GQA for the first time, in which MHA is converted to GQA using mean pooling initialization. However, this method requires uptraining to restore performance and incurs significant computational costs. Yu et al. (2024) keeps the corresponding parameters based on the principal components of the collected KV Caches, then uses LoRA Hu et al. (2021) to fine-tune the model to restore performance. Chen et al. (2024) proposes to regroup attention heads based on the criterion of cosine similarity and allows for varying group sizes. However, none of the aforementioned improvement methods can be fully adapted to the standard GQA model.
@@ 2.3 Compressing model based on the principal components of features
Some previous works Liu et al. (2023); Yu and Wu (2023) have pointed out that the features of LLMs are generally low-rank. Therefore, identifying and deleting the low-rank components of the model is an effective method for model compression.
@@ 3 Method
In this section, we will specifically describe our method. Our method consists of two parts, transformation of the network and pruning training. Transformation of the network represents applying orthogonal transformations to the projection matrices in order to increase the similarity between attention heads of the same group, so that we can increase efficiency of model optimization. The pruning training process combines pruning using $\mathit{L_{0}}$ masks Louizos et al. (2017) and knowledge distillation Gou et al. (2021).
@@ 3.1 Motivation
where $i,j$ are any two of attention heads in the same block, $n$ represents the $n^{th}$ token in this cache. As shown in Figure 2, we notice that while a few pairs of KV Caches share high cosine similarity with each other, the vast majority of them are almost orthogonal. This is the reason why directly mean-pooling projection matrices results in significant loss so that uptraining is needed to restore performance.
However, according to Yu et al. (2024), KV Caches are Low-rank. Given that these caches occupy only a portion of spatial dimensions, applying appropriate orthogonal transformations to the projection matrices to align key and value caches can reduce the difficulty of model optimization. Fortunately, this approach is feasible.
@@ 3.2 Preliminaries
Figure 3: This figure shows the average cosine similarity of key and value caches between any two heads before and after an applying transformation in some blocks of LLaMA2-7B. Appropriate orthogonal transformations can significantly improve the cosine similarity between KV Caches.
@@ 3.3 Transformation of attention heads
To calculate the optimal orthogonal matrix for each pair of key and value heads, we collect KV Caches according to the method mentioned above. Here, we use two criteria to perform the calculations.
@@
where the projection matrices in the attention heads are $W_{Q_{i}},W_{K_{i}},W_{V_{i}}\in\mathbb{R}^{d_{H}\times d}$ and $W_{O_{i}}\in\mathbb{R}^{d\times d_{H}}$, $X\in\mathbb{R}^{d\times len}$ represents the input sequence, and $x\in\mathbb{R}^{d\times 1}$ represents the current token. For brevity, RoPE is ignored here. Then we can fuse the orthogonal matrix into the value projection matrix ${W}_{V_{j}}$ and the output projection matrix $W_{O_{j}}$ to ensure computational invariance:
@@
As for $W_{Q}$ and $W_{K}$, due to the existence of RoPE, orthogonal transformation cannot be applied directly. However, we can divide the d-dimension space into $d/2$ sub-spaces and apply the orthogonal transformation in every two dimension just like RoPE, which is to say the orthogonal matrix for key projection matrix should be in this form:
@@
In the next section, we will compare the performances of the two criteria.
@@ 3.4 Find better grouping method
After obtaining the similarity scores between every pair of attention heads, we can regroup attention heads based on these scores. We define the similarity score of a group as the sum of similarity scores between every pair of attention heads within that group, and the total similarity score for each grouping result is the sum of similarity scores of all groups. Our objective is to determine the grouping result with the highest total score [^1] While the highest similarity between pairs within the same group does not necessarily equate to the lowest cost in terms of converging to the same parameters during pruning, this strategy remains acceptable considering computation and time costs.. We use $SimK^{after}$ and $SimV^{after}$ as grouping criteria, respectively. In the next section, we will compare the performances of these two ways. The mathematical expression of the score of a grouping result $A=\{A_{1},A_{2},\cdots,A_{G}\}$ is as follows:
@@
We use Simulated Annealing Algorithm to get the best grouping result: Exchange two random heads in different groups and calculate the new score, accepting the new result if it reaches a higher score. Repeat this process for multiple iterations. Because initialization has a significant impact on the final result, we run the algorithm multiple times. The details of the algorithm 2 are shown below.
@@
  Set $score_{best}$ $\leftarrow$-$\infty$ - ∞
@@
     $score_{current}$leftarrow calculate_score($G_{n}$, $SimV$)
@@
        $G^{{}^{\prime}}_{n}$leftarrow exchange random elements from different groups in $G_{n}$
        $score_{new}$leftarrow calculate_score($G^{{}^{\prime}}_{n}$, $SimV$)
@@
3.5 Adaptation of $\mathit{L_{0}}$ regularization
During pruning training, we add new projection matrices which are initialized by mean-pooling all the original heads within that group to the model Ainslie et al. (2023), here we use $\tilde{W}_{K_{k,g}}$ or $\tilde{W}_{V_{k,g}}$ to represent new projection matrices of the $g^{th}$ group in the $k^{th}$ block of the model:
$\tilde{W}_{K_{k,g}}=\frac{1}{D}\sum^{D}_{i=1}W_{K_{k,(g-1)*D+i}}$ (25)
$\tilde{W}_{V_{k,g}}=\frac{1}{D}\sum^{D}_{i=1}W_{V_{k,(g-1)*D+i}}$ (26)
These new matrices will be trained together with the model and replace original key-value heads after pruning. Assume the model has $N_{blocks}$ blocks and $H$ heads in an attention layer, we introduce $\mathit{L_{0}}$ masks $z\in\mathbb{R}^{N_{blocks}\times H}$ Louizos et al. (2017) to achieve this goal:
$W_{K_{k,j}}^{apply}=z_{k,j}W_{K_{k,j}}+(1-z_{k,j})\tilde{W}_{K_{k,g}}$ (27)
$W_{V_{k,j}}^{apply}=z_{k,j}W_{V_{k,j}}+(1-z_{k,j})\tilde{W}_{V_{k,g}}$ (28)
where $g=\lceil\frac{j}{D}\rceil$, $W_{K_{k,j}}$ and $W_{V_{k,j}}$ are the original projection matrices, $\tilde{W}_{K_{k,g}}$ and $\tilde{W}_{V_{k,g}}$ are the newly added projection matrices, $W_{K_{k,j}}^{apply}$ and $W_{V_{k,j}}^{apply}$ are the projection matrices employed during pruning. Following the $\mathit{L_{0}}$ regularization approach, we parametrize the pruning masks to hard concrete distributions. Initially, each mask is set $z=1$, we constrain the masks to zero during pruning Xia et al. (2023). And the original projection matrix will be transferred to the new matrix when $z=0$. Unlike traditional $\mathit{L_{0}}$ regularization, we aim to eliminate any original key or value heads and just utilize $\mathit{L_{0}}$ masks to gradually transfer the original heads to newly added heads. All masks across blocks are constrained by a single loss function:
~$\tilde{\mathcal{L}}_{L_{0}}=\left((\frac{1}{N_{block}H}\sum z)-T\right)+\left((\frac{1}{N_{block}H}\sum z)-T\right)^{2}$ (29)
We use vanilla KL loss and BiLD loss Li et al. (2024) to encourage alignment of student logits with teacher logits.
To sum up, the overall training loss is:
=~$\mathcal{L}=\tilde{\mathcal{L}}_{distill}+\tilde{\mathcal{L}}_{L_{0}}$ (31)
4 Experiments
4.1 Settings
Model configurations. We apply our method to the LLaMA2-7B model Touvron et al. (2023) throughout all experiments. We will convert the source model to GQA-16, GQA-8, and GQA-4, and compare them to the fine-tuned full-size model separately.
Datasets. We use the following open-source datasets for pruning training and evaluation: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), ARC-challengeClark et al. (2018), SIQA Sap et al. (2019) and OpenbookQA Mihaylov et al. (2018). The size and instruction template of each dataset are listed in Appendix B.
Implementation Details. We use 1 NVIDIA A100 GPU to perform model transformation, and 8 NVIDIA A100 GPUs for supervised fine-tuning (SFT) the teacher model and pruning training. We randomly select 128 sequences of 2048 tokens long from the C4 training set Raffel et al. (2020) as calibration data in model transformation. In all experiments, the initial learning rate is 1e-5 for the model parameters and 1e-2 for the pruning masks. The cosine scheduler is employed to reduce the learning rate to 0 by the end of training. We perform 2 epochs of SFT on the teacher model, 5 epochs of pruning training on GQA-16, 15 epochs on GQA-8, and 20 epochs on GQA-4.
More hyperparameter settings can be found in Appendix A.
4.2 Ablation studies
We test the impact of different similarity evaluation criteria (see Section 3.3) and grouping strategies (see Section 3.4). All results are presented in Table 1. Here, "baseline" refers to pruning directly without any transformation, "default grouping" refers to merging adjacent attention heads, "grouping by key" and "grouping by value" indicate grouping attention heads based on key or value cache similarity. "cos" and "dist" represent the transformation based on cosine similarity or Euclidean distance.
4.3 Main Results
We report the experimental results in Table 1. Except for one set of experiments, all transformed models outperform the baseline. As the sparsity of key-value head increases, the advantage of model transformation becomes more obvious, demonstrating the effectiveness of aligning the attention heads. While the results of the GQA-8 don’t meet expectations, in other experiments, models transformed based on value cache similarity, using Euclidean distance as the criterion, achieved the best performance.
@@ 4.3 Main Results
Table 1: Performances of prompt tuning on LLaMA2 models with various Methods. The last column, Avg. (Average Accuracy), indicates the average accuracy of all these sub datasets.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated exceptional performance across diverse natural language processing tasks. However, as the model size and the input sequence’s length increase, the linearly increasing key-value (KV) cache significantly degrades inference throughput. Therefore, grouped-query attention (GQA), as an alternative to multi-head attention (MHA), has been widely introduced into LLMs. In this work, we propose a cost-effective method for converting MHA into GQA with any compression ratio of KV heads. The key point of our method lies in the application of Procrustes analysis to the attention heads, which enhances the similarity among attention heads while preserving computational invariance, thereby improving the model’s post-training performance. Subsequently, we employ $\mathit{L_{0}}$ regularization to prune redundant parameters. The model after pruning can be adapted to the standard GQA framework. Experimental results show that our strategy can compress up to 87.5% KV heads of LLaMA2-7B model and 75% KV heads of Sheared-LLaMA-1.3B with acceptable performance degradation. Our code is released at https://github.com/fpcsong/mha2gqa.
@@ 1 Introduction
Recently, large language models (LLMs) Radford et al. (2018); Brown et al. (2020); Ouyang et al. (2022) show remarkable performance on a variety of natural language processing tasks. However, since most LLMs are based on Transformer architecture Vaswani et al. (2017), the expansion of the sequence length during inference leads to a linear increase in the memory footprint of key-value (KV) cache, which substantially increases on-device memory consumption. A reduced KV cache footprint not only lowers inference costs but also facilitates the processing of longer sequences and improves inference speed. Therefore, reducing the size of the KV cache is a key issue for LLMs.
Many methods for KV cache compression have been proposed, including KV cache quantization Hooper et al. (2024); Yue et al. (2024); Yang et al. (2024b), tokens dropping Adnan et al. (2024); Liu et al. (2023a); Tang et al. (2024) and so on. However, these approaches often introduce additional computational procedures, which are incompatible with general LLM frameworks.
@@
Another approach to KV cache compression is to directly change the attention architecture. Multi-query attention (MQA) Shazeer (2019) and grouped-query attention (GQA) Ainslie et al. (2023) reduce KV cache by allowing multiple attention heads to share a single KV head, offering a simple and effective solution for attention optimization. Since GQA has better inference stability and performance, it has been widely used in LLaMA 2 Touvron et al. (2023), LLaMA 3 Dubey et al. (2024), Qwen2 Yang et al. (2024a), Mistral Jiang et al. (2023) and other LLMs Liu et al. (2024c); Zhang et al. (2024). Multi-head latent attention (MLA) Liu et al. (2024a) further reduces KV cache through low-rank projection of the cached data. MLA has been used in the DeepSeek model series Liu et al. (2024a, b). These efficient attention architectures achieve KV cache compression with better universality.
In this study, we present a method that converts MHA to GQA to compress KV cache. Inspired by the idea of computational invariance in LLMs Ashkboos et al. (2024) and Procrustes analysis Schönemann (1966), we apply proper orthogonal transformations to the projection matrices in attention heads to simplify the conversion from MHA to GQA: Specifically, we regroup attention heads based on the similarity of their KV caches and use generalized Procrustes analysis Wikipedia contributors (2022) to maximize the similarity of attention heads within each group. This transformation preserves the model’s output invariance. Finally, $\mathit{L_{0}}$ regularization Louizos et al. (2017) is applied to transfer original KV heads to new ones to get GQA. Figure 1 illustrates the overall framework of our method. Experimental results show that attention head transformation can significantly improve the performance of the pruned model. Our contributions are as follows.
  • Based on the idea of computational invariance, we employ Procrustes analysis to enhance the similarity among attention heads. This approach not only improves the performance of GQA model, but also provides a new perspective for evaluating the similarity between attention heads, offering new insights for future research related to compressing KV cache.
@@
  • We propose a general and cost-effective method for converting MHA to GQA, using $\mathit{L_{0}}$ regularization to compress the key and value heads to any percentage and basically restore performance after supervised fine-tuning.
@@
  • We conduct experiments on LLaMA2-7B Touvron et al. (2023) and Sheared-LLaMA-1.3B Xia et al. (2023), converting them into GQA models of varying sizes, separately. Model performance does not decrease significantly compared to that of MHA model.
    Figure 1: An illustration of our method. 1. In each attention layer of LLM, after the average similarity score is evaluated by Procrustes analysis between every two value caches (or key caches). Use Simulated Annealing Algorithm to search the optimal grouping result that maximizes the similarity score among value caches (or key caches) in each layer. 2. After grouping, then fuse the orthogonal matrix obtained from generalized Procrustes analysis into each projection matrix to enhance the similarity among attention heads in each group without changing the model. 3. During training, $\mathit{L_{0}}$ loss is used to gradually transfer original key-value projection matrices to newly added ones within each group (there are $G$ groups in one attention layer). After pruning, original key-value projection matrices will be discarded, then we get a standard GQA model.
@@ 2.2 Transfer MHA to GQA
Ainslie et al. (2023) proposes GQA for the first time, in which MHA is converted to GQA using mean pooling initialization. However, this method requires uptraining to restore performance and incurs significant computational costs. Yu et al. (2024) keeps the corresponding parameters based on the principal components of the collected KV caches, then uses LoRA Hu et al. (2021) to fine-tune the model to restore performance. Chen et al. (2024) proposes to regroup attention heads based on the criterion of cosine similarity and allows for varying group sizes. DHAChen et al. (2024a) adaptively configures group sharing for key heads and value heads across various layers, achieving a better balance between performance and efficiency. However, none of the aforementioned improvement methods can be fully adapted to the standard GQA model.
@@ 2.3 Compressing model based on the principal components of features
Some previous works Liu et al. (2023b); Yu and Wu (2023) have pointed out that the features of LLMs are generally low-rank. Therefore, identifying and deleting the low-rank components of the model is an effective method for model compression.
@@ 3 Method
In this section, we will specifically describe our method. Our method consists of two parts, transformation of attention heads and pruning training. Transformation of attention heads represents employing Procrustes analysis to align projection matrices in order to enhance the similarity between attention heads of the same group, so that we can increase efficiency of merging attention heads. The pruning training process combines pruning with $\mathit{L_{0}}$ regularization Louizos et al. (2017) and knowledge distillation Gou et al. (2021).
@@ 3.1 Motivation
where $i,j$ are any two of attention heads in the same block, $n$ represents the $n^{th}$ token in this cache. Taking LLaMA2-7B as an example, as shown in Figure 2, we notice that the vast majority of KV caches are almost orthogonal.
However, according to Yu et al. (2024), KV caches are low-rank. Given that these caches occupy only a portion of spatial dimensions, applying appropriate orthogonal transformations to the projection matrices to align key and value caches can increase efficiency of merging attention heads. Fortunately, this approach is feasible.
@@ 3.2 Preliminaries
Figure 2: This figure shows the average cosine similarity of key and value caches between any two heads before and after applying transformation in some blocks of LLaMA2-7B. It can significantly improve the cosine similarity between KV caches after Procrustes analysis (In fact, it can also significantly reduce the Euclidean distance between any two caches.).
@@ 3.3 Transformation of attention heads
To calculate the optimal orthogonal matrix for each pair of key and value heads, we collect KV caches according to the method mentioned above. Here, we use two criteria to perform Procrustes analysis.
@@
where the projection matrices in the attention heads are $W_{Q_{i}},W_{K_{i}},W_{V_{i}}\in\mathbb{R}^{d_{H}\times d}$ and $W_{O_{i}}\in\mathbb{R}^{d\times d_{H}}$, $X\in\mathbb{R}^{d\times len}$ represents the previous tokens, and $x\in\mathbb{R}^{d\times 1}$ represents the current token. For brevity, Rotary position embedding (RoPE) is ignored here. Then we can fuse the orthogonal matrix into the value projection matrix ${W}_{V_{j}}$ and the output projection matrix $W_{O_{j}}$ to ensure computational invariance:
@@
As for $W_{Q}$ and $W_{K}$, due to the existence of RoPE Su et al. (2024), Procrustes analysis cannot be applied directly. However, we can divide the d-dimension space into $d/2$ sub-spaces and apply Procrustes analysis in every two dimension just like RoPE, which is to say the orthogonal matrix for key projection matrix should be in this form:
@@
In the next section, we will compare the performance of the two criteria.
@@ 3.4 Find better grouping method
After obtaining the similarity scores between every pair of attention heads, we can regroup attention heads based on these scores. We define the similarity score of a group as the sum of similarity scores between every pair of attention heads within that group, and the total similarity score for each grouping result is the sum of similarity scores of all groups. Our objective is to identify the grouping result with the highest total score [^1] While the highest similarity between pairs within the same group does not necessarily equate to the lowest cost in terms of converging to the same parameters during pruning, this strategy remains acceptable when considering computation and time costs.. We use $SimK^{after}$ and $SimV^{after}$ as grouping criteria, respectively. In the next section, we will compare the performance of the two criteria. The mathematical expression of the score of a grouping result $A=\{A_{1},A_{2},\cdots,A_{G}\}$ is as follows:
@@
We use Simulated Annealing Algorithm to get the best grouping result: Swap two random heads in different groups and calculate the new score, accepting the new result if it reaches a higher score. Repeat this process for multiple iterations. Because initialization has a significant impact on the final result, we run the algorithm multiple times. The details of the algorithm 2 are shown below.
@@
 Set $score_{best}$leftarrow-$\infty$ - ∞
@@
[…]
@@
   $G^{{}^{\prime}}_{n}$leftarrow Randomly swap two elements from different groups in $G_{n}$
[…]
@@
Table 1: Performances of LLaMA2-7B with various methods.
3.5 Adaptation of $\mathit{L_{0}}$ regularization
During pruning training, we add new projection matrices which are initialized by mean-pooling all the original heads within that group to the model Ainslie et al. (2023), here we use $\tilde{W}_{K_{b,g}}$ or $\tilde{W}_{V_{b,g}}$ to represent new projection matrices of the $g^{th}$ group in the $b^{th}$ block of the model:
$\tilde{W}_{K_{b,g}}=\frac{1}{D}\sum^{D}_{i=1}W_{K_{b,(g-1)*D+i}}$ (25)
$\tilde{W}_{V_{b,g}}=\frac{1}{D}\sum^{D}_{i=1}W_{V_{b,(g-1)*D+i}}$ (26)
These new matrices will be trained together with the model and replace original KV heads after pruning. Assume the model has $N_{blocks}$ blocks and $H$ heads in an attention layer, we introduce $\mathit{L_{0}}$ masks $z\in\mathbb{R}^{N_{blocks}\times H}$ Louizos et al. (2017) to achieve this goal:
$W_{K_{b,j}}^{apply}=z_{b,j}W_{K_{b,j}}+(1-z_{b,j})\tilde{W}_{K_{b,g}}$ (27)
$W_{V_{b,j}}^{apply}=z_{b,j}W_{V_{b,j}}+(1-z_{b,j})\tilde{W}_{V_{b,g}}$ (28)
where $g=\lceil\frac{j}{D}\rceil$, $W_{K_{b,j}}$ and $W_{V_{b,j}}$ are the original projection matrices, $\tilde{W}_{K_{b,g}}$ and $\tilde{W}_{V_{b,g}}$ are the newly added projection matrices, $W_{K_{b,j}}^{apply}$ and $W_{V_{b,j}}^{apply}$ are the actual projection matrices employed during pruning. Following the $\mathit{L_{0}}$ regularization approach, we parametrize the pruning masks to hard concrete distributions. Initially, each mask is set $z=1$, we constrain the masks to zero during pruning Xia et al. (2023). And the original projection matrix will be transferred to the new matrix when $z=0$. Unlike traditional $\mathit{L_{0}}$ regularization, we aim to eliminate any original key or value heads and just utilize $\mathit{L_{0}}$ masks to gradually transfer the original heads to newly added heads. All masks across blocks are constrained by a single loss function:
~$\tilde{\mathcal{L}}_{L_{0}}=|(\frac{1}{N_{block}H}\sum z)-T|+\left((\frac{1}{N_{block}H}\sum z)-T\right)^{2}$ (29)
Before pruning, we already SFT an MHA model as teacher model. Then we use vanilla KL loss and BiLD loss Li et al. (2024) to encourage alignment of student logits with teacher logits.
To sum up, the overall training loss is:
=~$\mathcal{L}=\tilde{\mathcal{L}}_{distill}+\tilde{\mathcal{L}}_{L_{0}}$ (31)
4 Experiments
4.1 Settings
Model configurations. We apply our method to the LLaMA2-7B model Touvron et al. (2023) and Sheared-LLaMA-1.3B Xia et al. (2023) throughout all experiments. We will prune the KV heads of MHA at different pruning rates, and compare them to the fine-tuned MHA model separately.
Datasets. We use the following open-source datasets for training and evaluation: BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-easy Clark et al. (2018), ARC-challengeClark et al. (2018), SIQA Sap et al. (2019) and OpenbookQA Mihaylov et al. (2018). The size and instruction template of each dataset are listed in Appendix B.
Implementation Details. We use 1 A100 GPU to perform model transformation, and 8 A100 GPUs for SFT the teacher model and pruning training. We randomly select 128 sequences of 2048 tokens long from the C4 training set Raffel et al. (2020) as calibration data in attention heads transformation. During the transformation, we convert the model parameters to float64 to reduce the calculation error. The time required for calibration and conversion is detailed in Appendix C. In pruning training, the initial learning rate is 1e-5 for the model parameters and 1e-2 for the pruning masks. The cosine scheduler is employed to reduce the learning rate to 0 by the end of training. More hyperparameter settings can be found in Appendix A.
Budgets
(tokens and epochs)
MHA Teacher 84.83 72.74 68.31 71.35 53.18 64.04 59.20 71.34 71.37
55.8M
3 epochs
GQA-8 baseline 81.99 69.10 58.22 71.51 44.82 61.05 60.20 68.73 64.99
111.6M
6 epochs
default grouping cos 83.27 70.95 63.60 71.43 50.50 62.63 58.60 70.52 68.38
dist 83.85 71.11 63.54 71.35 50.50 63.68 58.60 71.03 68.54
grouping by key cos 82.72 69.75 62.42 70.96 46.49 61.58 59.60 68.68 67.29
dist 82.11 65.34 62.61 69.38 40.80 54.21 49.80 66.22 65.99
grouping by value cos 83.06 70.35 63.63 72.14 49.83 61.75 60.40 70.32 68.33
dist 83.55 71.21 63.54 73.56 46.82 61.75 60.60 70.57 68.53
GQA-4 baseline 79.33 64.53 50.42 70.64 38.13 57.02 54.80 66.79 59.55
223.2M
12 epochs
default grouping cos 81.41 68.93 59.30 71.98 42.14 57.54 56.00 69.04 65.24
dist 81.10 69.20 59.47 70.80 43.81 59.65 57.20 68.89 65.33
grouping by key cos 81.99 68.23 59.94 71.43 45.15 58.60 59.00 68.83 65.69
dist 81.50 67.95 59.54 71.35 47.83 57.54 59.80 67.96 65.32
grouping by value cos 81.25 68.28 60.06 70.32 47.16 60.18 60.40 69.14 65.71
dist 82.29 67.63 59.77 70.40 46.82 60.00 60.20 69.14 65.66
Table 2: Performances of Sheared-LLaMA-1.3B with various methods.
4.2 Ablation studies
We test the impact of different similarity evaluation criteria (see Section 3.3) and grouping strategies (see Section 3.4). All results are presented in Table 1 and Table 2. Here, "baseline" refers to pruning directly without any transformation, "default grouping" refers to merging adjacent attention heads, "grouping by key" and "grouping by value" indicate grouping attention heads based on key or value cache similarity. "cos" and "dist" represent the transformation based on cosine similarity or Euclidean distance.
4.3 Main results
We report the experimental results and budgets in Table 1 and Table 2, Avg. (Average Accuracy), indicates the average accuracy of all these sub datasets. Except for one set of experiments, all transformed models outperform the baseline. As the sparsity of KV head increases, the advantage of model transformation becomes more obvious, demonstrating the effectiveness of aligning attention heads.